OpenStudy (anonymous):

Find the modulus and argument of each root: i+2 and i-2

OpenStudy (shubhamsrg):

lol..is that the continuation of your previous ques ?

OpenStudy (anonymous):

Yup

OpenStudy (shubhamsrg):

well..i'll explain geometrically.. modulus of a complex number x+iy means the distance of (x,y) from (0,0) which is => (x^2 + y^2)^1/2 and argument is the angle which the line joining point (x,y) and (0,0) make with positive x axis.. i hope you should again be able to do now..

OpenStudy (anonymous):

Still not sure. Which is x and which is y?

OpenStudy (shubhamsrg):

coeff of i is y and the other one is x so in i+2 ,, x=2 and y=1 in i-2 ,, x=-2 and y =1

OpenStudy (anonymous):

So \[\sqrt 5\]

OpenStudy (anonymous):

and what of argument?

OpenStudy (shubhamsrg):

have you learnt about slopes in 2-d geometry ?

OpenStudy (anonymous):

Yes, I have.

OpenStudy (shubhamsrg):

then you must be able to calculate slope of line joining (x,y) and (0,0) .. the slope means tan theta..which gives you theta..which is what you want..

OpenStudy (anonymous):

I tried for the second one, but I didn't get the correct value...I did for the first though.

OpenStudy (anonymous):

I got it. Don't worry.

OpenStudy (anonymous):

The last part is to draw an argand diagram, however. How?

OpenStudy (anonymous):

Don't worry/

OpenStudy (shubhamsrg):

you want me to draw it ? or you got it?

OpenStudy (anonymous):

I got it

OpenStudy (shubhamsrg):

great then miss emma watson :P !!

OpenStudy (anonymous):

how to go about finding the argument for \[(1/2) + i (\sqrt{3}/2)\] ??

OpenStudy (shubhamsrg):

calculate the slope of line joining (1/2 , sqrt(3) /2) and (0,0) slope = tan theta thus find theta which is the argument..